OpenStudy (dilla):

a boat can travel 140 miles downstream in the same time that it takes to travel 92 miles upstream. The speed of the current is 6mph. What is the speed of the boat in still water?

OpenStudy (dilla):

@triciaal

OpenStudy (triciaal):

@mathmate please help I'm about to leave

OpenStudy (mathmate):

@triciaal Sorry, I wasn't online. Let the boat speed in stillwater be S, and speed of current be C. Then S+C=downstream speed S-C=upstream speed Use Distance = Speed * Time => time = Distance / Speed We are given time to travel 140 miles downstream is the same as time to travel 90 miles upstream, which makes 140/(S+C)=90/(S-C) Since we know the current speed C=6, above simplifies to 140/(S+6)=90/(S-6) Cross multiply: 140(S-6)=90(S+6) solve for S.